# InternVLA-A1: Unifying Understanding, Generation and Action for Robotic Manipulation

### v1

@@ 3.3 Optimization Objectives
Our training process consists of two sequential stages: Pre-training and Post-training. Although these two stages utilize distinct data sources and training hyperparameters (detailed in Table 2), they share a unified optimization framework and identical objectives. Throughout both stages, we jointly optimize the model for two key objectives: visual foresight generation and action prediction.
(1) Visual Foresight Generation. To endow the model with predictive capabilities about future visual states, we train the generation expert to forecast the latent representation of a future frame. Let denote the Cosmos latent. Given the current and historical observations and language instruction , the generation expert predicts conditioned on the understanding prefix and is supervised by the frozen Cosmos latent . This objective is formulated as:
where denotes the generation expert, is the frozen Cosmos encoder used in , and indicates the stop-gradient operation. This objective compels the model to internalize physical dynamics, creating a robust prior that enhances action prediction.
(2) Flow Matching-based Action Prediction. To establish a principled connection between visual imagination and motor control, we employ a flow matching framework for action learning. This approach parameterizes the action generation process as learning continuous transformation pathways from noise to expert demonstrations, offering superior handling of multimodal action distributions compared to direct regression. Formally, we sample time steps and construct interpolated action chunks , where represents Gaussian noise. The model learns a velocity field that transports noisy samples toward target actions:
where denotes the proprioception state at time . During inference, sampling from the learned policy distribution is achieved by solving an ODE: starting from Gaussian noise , we iteratively apply the Euler update:
@@ 3.4 Implementation Details
Regarding inference efficiency, both InternVLA-A1 (2B) and InternVLA-A1 (3B) run at approximately 13 Hz with torch.compile. Notably, InternVLA-A1 (2B) does not exhibit lower latency than InternVLA-A1 (3B) despite having fewer total parameters. This is because the InternVL3 backbone in InternVLA-A1 (2B) requires a higher input resolution () compared to the input used by the Qwen3-VL backbone in InternVLA-A1 (3B). Consequently, the computational cost of processing longer visual token sequences in InternVLA-A1 (2B) offsets its parameter advantage, leading to comparable overall speeds.
@@
Training Hyperparameters. We train InternVLA-A1 in two stages: a large-scale pre-training stage on heterogeneous robotic datasets, followed by a task-specific post-training stage. During pre-training, we optimize the model using AdamW with a constant learning rate schedule for 700K steps. For post-training, we adopt a lower learning rate with a warmup and decay schedule to stabilize adaptation to downstream tasks. All experiments are conducted using bfloat16 precision. The hyperparameter , which balances the two loss components, is set to 0.01. The interval between the historical frame and the current frame, as well as between the future frame and the current frame, is set to 15.
@@
Load-balanced Parallel Training. InternVLA-A1 is trained on a mixture of heterogeneous robot datasets. At this scale, naively instantiating the full dataset mixture on every worker can trigger out-of-memory failures and exacerbate I/O contention. We therefore adopt Load-balanced Parallel Training (LPT), a distributed data-loading strategy that assigns datasets to workers to achieve both scalability and statistically well-behaved sampling.
@@ 4 Data Corpus
4.1 Hierarchical data pyramid
This study proposes a hierarchical data pyramid framework, as presented in Figure˜3. This data pyramid strategically integrates diverse data sources across training stages. The foundation is built during pre-training, where we utilize large-scale open-source robot demonstrations alongside high-fidelity synthetic data to establish broad generalizability. This is followed by a post-training phase that leverages specialized real-world data, refining the model’s capabilities for precise physical deployment.
@@ 4.2 Simulated synthetic data
We incorporate our prior work, InternData-A1 (tian2025interndata_a1), a large-scale synthetic robot dataset that is among the most diverse and comprehensive to date. The dataset contains over 630k trajectories and 7,433 hours of data spanning 4 embodiments, 18 skills, 70 tasks, and 227 scenes, covering manipulation of rigid, articulated, deformable, and fluid objects. It is generated via a highly autonomous, fully decoupled, and compositional simulation pipeline, enabling long-horizon skill composition, flexible task assembly, and support for heterogeneous embodiments with minimal manual tuning. InternData-A1 is the first to demonstrate that synthetic-only data can match the performance of large-scale real-world datasets when pre-training VLA models, achieving comparable results to the strongest closed-source real-world -dataset (black2024pi_0). Moreover, models trained on InternData-A1 exhibit strong zero-shot sim-to-real transfer on several challenging tasks. As the foundation of our data corpus, InternData-A1 provides rich diversity in trajectories, objects, and environments. We select InternData-A1 as the foundation of our pre-training corpus due to its exceptional sample diversity and proven efficacy in pre-training VLA models.
4.3 Open-source robot demonstration
@@
While synthetic data excels in scalability, real-world demonstrations remain essential for capturing authentic physical dynamics and bridging the sim-to-real gap. When selecting real-world demonstration data for our pre-training corpus, we prioritize datasets that exhibit large-scale trajectory coverage, diverse task distributions, and high-quality teleoperation. Based on these criteria, we incorporate the open-source AgiBot-World dataset (bu2025agibotworld). This dataset contains over 1 million trajectories covering 217 tasks across five domains, including domestic, retail, industrial, restaurant, and office scenarios. By incorporating this large-scale real-world demonstration dataset into our pre-training corpus, we benefit from its high-quality and diverse demonstrations, which complement our synthetic data and help bridge the sim-to-real gap.
4.4 Specialized real-world data
At the apex of our data pyramid, we utilize specialized real-world data to fine-tune the model for specific downstream applications. Collected via teleoperation in targeted environments, these high-quality, task-centric episodes—though smaller in scale—are highly aligned with the target domain. Post-training on this data effectively adapts the model’s broad pre-trained knowledge to the precise nuances of the deployment environment.
4.5 Pre-training data recipe
We pretrain the model on a mixture of large-scale simulated synthetic data and open-source real-world demonstration data. The synthetic portion is drawn from InternData-A1, which covers four robot embodiments (Genie-1, Franka, ARX Lift-2, and AgileX) with diverse task families and long-horizon compositions. We additionally include AgiBot-World as real-world data to complement synthetic trajectories with physical interaction variability. During pretraining, we interleave trajectories from different sources using configurable sampling weights (Table 3).
| Data source | Num. frame | Frame proportion |
| InternData-A1 (ARX Lift-2) | 96M | 0.18 |
| InternData-A1 (AgileX) | 122.5M | 0.23 |
| InternData-A1 (Franka) | 90.5M | 0.17 |
| InternData-A1 (Genie-1) | 16M | 0.03 |
| AgiBot-World (Beta) | 208M | 0.39 |
@@ 5 Experiments
To evaluate the effectiveness of our proposed InternVLA-A1, we conduct extensive experiments on twelve real-world tasks and a simulation benchmark. We first compare InternVLA-A1 with existing leading vision-language-action models and GR00T N1.5. Then, we conduct an ablation study to verify the effectiveness of pretraining and the proposed generation expert.
@@ 5.1 Evaluation Configuration
We benchmark our policy on three physical robot embodiments: Genie-1, ARX Lift-2, and ARX AC One. These platforms cover diverse bimanual manipulation capabilities and execution characteristics, enabling us to assess real-robot performance across both long-horizon and contact-rich behaviors under consistent sensing and control setups. All evaluations are performed on real hardware with the same deployment pipeline across the three robots, and we report task success rates based on repeated rollouts.
@@
Task setup. We design a real-world task suite that covers both general-purpose manipulation and specialized high-dynamics scenarios. The suite includes ten general-purpose tasks and two specialized task families:
General-purpose tasks:
Dexterous manipulation: Unscrew Cap, Zip Bag, Sort Parts.
@@
Regular manipulation: Operate Oven, Make Sandwich, Sort Rubbish, Sweep Trash, Wipe Stain, Place Markpen, Place Flower.
@@
Specialized tasks (high dynamics): In-motion Ingredient Picking tasks and Express Sorting tasks.
@@
Overall, the tasks span articulated-object interaction and contact-rich manipulation (e.g., cap removal and zipper closure), long-horizon bimanual routines (e.g., sweeping/disposing and multi-object sorting), and, importantly, high-dynamics settings where the scene evolves during execution. The latter requires the policy to reason about near-future scene changes, so that actions can be generated with upcoming dynamics in mind rather than solely based on the current observation.
@@
Evaluation protocol. We report the average success rate (and grasp success when applicable) over 30 rollouts per task. Specifically, each task is evaluated under 30 predefined settings (e.g., object placements and scene initializations within bounded ranges), with one trial per setting. Results are averaged across all rollouts to summarize performance across diverse initial conditions.
5.2 Evaluation on the General-purpose Tasks
@@
To evaluate the performance of InternVLA-A1 against leading VLA models, including and GR00T N1.5, we conducted a comprehensive evaluation across 10 diverse real-world tasks. These tasks, as displayed in Figure˜4, range from dexterous manipulation (e.g., Unscrew Cap, Zip Bag, Sort Parts) to regular manipulation (e.g., Make Sandwich, Operate Oven, Sort Rubbish).
@@
As shown in Table 4, InternVLA-A1 demonstrates superior performance compared to the prior state-of-the-art models, (3.3B) and GR00T N1.5 (3B). Notably, our smaller InternVLA-A1 (2B) model achieves an average success rate of 64.7%, surpassing the 60.6% average of the larger (3.3B) baseline. This result highlights the efficiency of our architecture and the quality of our data pyramid, enabling a lighter model to outperform a larger parameter counterpart. When scaling to a comparable size, InternVLA-A1 (3B) yields a substantial performance improvement, reaching an average success rate of 75.1%—a 14.5% absolute improvement over . In regular manipulation tasks, InternVLA-A1 (3B) exhibits exceptional robustness; it achieves dominant leads in complex tasks like Make Sandwich (93.3% vs. 66.7%) and Operate Oven (86.7% vs. 73.3%), while maintaining top-tier performance on Sort Rubbish (97.3%), matching the strong baseline set by . Furthermore, in challenging dexterous manipulation tasks, InternVLA-A1 (3B) equals the strongest baseline in Unscrew Cap and Sort Parts while significantly surpassing it in Zip Bag (73.3% vs. 40.0%). These results confirm that InternVLA-A1 not only excels in high-level task planning but also possesses superior fine-grained control capabilities.
@@
5.3 Evaluation on the Specialized Tasks
To further validate the performance of our proposed InternVLA-A1, we designed two tasks involving manipulation in dynamic scenes: Express Sorting and In-motion Ingredient Picking. The experimental settings for the Express Sorting and In-motion Ingredient Picking tasks are illustrated in Figure˜5. Both involve long-horizon operations within dynamic environments. In the Express Sorting Task, the robot must first determine if the package label is facing upwards. If the label is facing downwards, a four-step sequence is initiated wherein the right arm first executes a “chasing” grasp along the direction of the conveyor movement, followed by flipping the package, after which the left arm performs a “waiting” grasp. Finally, the package is lifted to present the label to the head-mounted camera. In the In-motion Ingredient Picking tasks, two robots coordinate to grasp the ingredients required to assemble a beef sandwich, consisting of two slices of bread, a steak, and a piece of lettuce.
As shown in Figure 6, the experimental results reveal that state-of-the-art baselines, and GR00T N1.5, struggle significantly in dynamic manipulation scenarios, yielding consistently low success rates. Specifically, these models fail to exceed 40.0% on the Express Sorting task and achieve only 20.0% on the In-motion Ingredient Picking task. In stark contrast, InternVLA-A1 (2B) and InternVLA-A1 (3B) demonstrate superior stability and robustness across both settings. The 2B model achieves success rates of 70.0% and 46.7%, respectively. Most notably, InternVLA-A1 (3B) exhibits exceptionally strong performance, attaining 80.0% in Express Sorting and a remarkable 93.3% in In-motion Ingredient Picking. This represents a substantial margin over the baselines, particularly in the multi-robot coordination task, where the 3B model outperforms and GR00T N1.5 by over 70 percentage points.
@@ 5.4 Evaluation on Simulation Benchmark
| Method | Hanging Mug | Move Pillbottle Pad | Pick Dual Bottles | Place Object Stand | |||||
| Easy | Hard | Easy | Hard | Easy | Hard | Easy | Hard | ||
| 9.0% | 5.0% | 30.0% | 0.0% | 58.0% | 16.0% | 47.0% | 14.0% | ||
| InternVLA-A1(2B) | 17.0% | 13.0% | 54.0% | 8.0% | 70.0% | 25.0% | 67.0% | 25.0% | |
| InternVLA-A1(3B) | 24.0% | 13.0% | 69.0% | 8.0% | 45.0% | 23.0% | 70.0% | 38.0% | |
| Blocks Ranking Size | Shake Bottle | Stamp Seal | Place Burger Fries | Avg. (50 Tasks) | |||||
| Easy | Hard | Easy | Hard | Easy | Hard | Easy | Hard | Easy | Hard |
| 19.0% | 1.0% | 96.0% | 53.0% | 16.0% | 3.0% | 66.0% | 34.0% | 54.5% | 19.8% |
| 27.0% | 7.0% | 99.0% | 64.0% | 42.0% | 6.0% | 99.0% | 10.0% | 58.6% | 20.5% |
| 26.0% | 8.0% | 99.0% | 79.0% | 39.0% | 12.0% | 96.0% | 52.0% | 65.0% | 25.4% |
We evaluate our method on the RoboTwin 2.0 benchmark (robotwin2.0), which comprises 50 bimanual tasks. The evaluation includes two protocols: the Easy setting, identical to the demonstration environment, and the Hard setting, featuring domain randomization. For each task, we fine-tuned the pre-trained model using 100 demonstrations and conducted 100 evaluation trials for both settings. Table 5 demonstrates that InternVLA-A1 surpasses the model across both the 2B and 3B variants. Specifically, the InternVLA-A1(3B) model yields performance gains of 10.5% in the easy setting and 5.6% in the hard setting.

### v2

@@ 3.3 Optimization Objectives
Our training process consists of two sequential stages: Pre-training and Post-training. Although these two stages utilize distinct data sources and training hyperparameters (detailed in Table 2), they share a unified optimization framework and identical objectives. Throughout both stages, we jointly optimize the model for two key objectives: visual foresight generation and action prediction. Let denote the training corpus for visual foresight (including human video data without action labels), and denote the subset with action annotations. We define and as the corresponding training tuples.
(1) Visual Foresight Generation. To endow the model with predictive capabilities about future visual states, we train the generation expert to forecast the latent representation of a future frame. Let denotes the COSMOS VAE encoder and denotes the COSMOS latent feature. Conditioned on current and historical observations , as well as the understanding prefix , the generation expert predicts the latent feature for the future timestamp . The prediction is supervised by the ground truth COSMOS latent feature . We minimize the following objective:
where denotes the generation expert, and indicates the stop-gradient operation. This objective compels the model to internalize physical dynamics, creating a robust prior for action.
(2) Flow Matching-based Action Prediction. We employ a flow matching framework for action learning. This approach formulates the action generation process as learning continuous transformation pathways from noise to expert demonstrations, offering superior handling of multi-modal action distributions compared to direct regression. Formally, given , we sample time steps and construct interpolated action chunks , where represents Gaussian noise. The model learns a velocity field that transports noisy samples toward target actions:
where denotes the proprioception state at time , and and are the contextual conditioning features produced by the understanding expert and the generation expert, respectively. During inference, sampling from the learned policy distribution is achieved by solving an ODE: starting from Gaussian noise , we iteratively apply the Euler update:
@@ 3.4 Implementation Details
Regarding inference efficiency, both InternVLA-A1 (2B) and InternVLA-A1 (3B) run at approximately 13 Hz with torch.compile on one RTX 4090 GPU. Notably, InternVLA-A1 (2B) does not exhibit lower latency than InternVLA-A1 (3B) despite having fewer total parameters. This is because the InternVL3 backbone in InternVLA-A1 (2B) requires a higher input resolution compared to the input used by the Qwen3-VL backbone in InternVLA-A1 (3B). Consequently, the computational cost of processing longer visual token sequences in InternVLA-A1 (2B) offsets its parameter advantage, leading to comparable overall speeds.
@@
Training Protocol and Hyperparameters. We train InternVLA-A1 in two stages: a large-scale pre-training stage on heterogeneous datasets, followed by a task-specific post-training stage. During pre-training, we optimize the model using AdamW with a constant learning rate schedule for 700K steps. For post-training, we adopt a lower learning rate with a warmup and decay schedule to stabilize adaptation to downstream tasks. Detailed hyperparameter settings for pre-training and post-training are provided Table˜2. In addition, the hyperparameter , which balances the two loss components, is set to 0.01. The interval between the historical frame and the current frame, as well as between the future frame and the current frame, is set to 15.
@@
Load-balanced Parallel Training. InternVLA-A1 is trained on a mixture of heterogeneous datasets over 692 frames. At this scale, naively instantiating the complete datasets on every worker like LeRobot codebase [cadene2024lerobot] can trigger out-of-memory issues and exacerbate I/O contention. We therefore adopt Load-balanced Parallel Training (LPT), a distributed data-loading strategy that assigns datasets to workers to achieve both scalability and statistically well-behaved sampling.
@@ 4 Data Corpus
4.1 Pre-training data recipe
We pretrain the model on a mixture of heterogeneous data sources including synthetic simulation data, real-world robot data, and human videos. The pre-training data recipe is shown in Table 3. During pretraining, we interleave trajectories from different sources using configurable sampling weights.
| Data source | Type | Num. frame | Sampling weight |
| InternData-A1 | Sim. | 396M | 0.64 |
| RoboTwin | Sim. | 17M | 0.08 |
| AgiBot-World (Beta) | Real | 206M | 0.18 |
| RoboMind | Real | 5M | 0.02 |
| EgoDex | Human | 68M | 0.08 |
@@ 4.2 Simulated synthetic data
We incorporate our prior work, InternData-A1 [tian2025interndata_a1], a large-scale synthetic robot dataset that is among the most diverse and comprehensive to date. The dataset contains over 630k trajectories and 7,433 hours of data spanning 4 embodiments, 18 skills, 70 tasks, and 227 scenes, covering manipulation of rigid, articulated, deformable, and fluid objects. It is generated via a highly autonomous, fully decoupled, and compositional simulation pipeline, enabling long-horizon skill composition, flexible task assembly, and support for heterogeneous embodiments with minimal manual tuning. InternData-A1 is the first to demonstrate that synthetic-only data can match the performance of large-scale real-world datasets when pre-training VLA models, achieving comparable results to the strongest closed-source real-world -dataset [black2024pi_0]. Moreover, models trained on InternData-A1 exhibit strong zero-shot sim-to-real transfer on several challenging tasks. With InternData-A1’s synthesis pipeline and system optimization enabled by the Nimbus [he2026nimbus], it generates 209.7 hours of simulation data per day on 8 RTX 4090 GPUs. As the foundation of our data corpus, InternData-A1 provides rich diversity in trajectories, objects, and environments. We select InternData-A1 as the foundation of our pre-training corpus due to its exceptional sample diversity and proven efficacy in pre-training VLA models. Additionally, we incorporate the simulation dataset proposed in RoboTwin [robotwin2.0].
4.3 Real-world robot data
@@
While synthetic data excels in scalability, real-world demonstrations remain essential for capturing nuanced physical dynamics and bridging the sim-to-real gap. When selecting real-world demonstration data for our pre-training corpus, we prioritize datasets that exhibit large-scale trajectory coverage, diverse task distributions, and high-quality teleoperation. Based on these criteria, we incorporate the open-source AgiBot-World dataset [bu2025agibotworld] and RoboMind [wu2024robomind]. By incorporating these large-scale real-world demonstration datasets into our pre-training corpus, we benefit from its high-quality and diverse demonstrations, which complement our synthetic data and help bridge the sim-to-real gap.
4.4 Egocentric human video
We additionally incorporate human video data into our data corpus. To align with robot manipulation scenarios, we prioritize egocentric videos that share a similar viewpoint and feature diverse interactions in robot-like environments. Specifically, we utilize the EgoDex [egodex] dataset, a large-scale collection focused on egocentric dexterous manipulation. It comprises 829 hours of footage spanning over 200 tasks. Notably, we exclude human action labels during pre-training. Rich in human-hand-object interactions, this data is instrumental for the generation expert, enabling it to capture the nuanced dynamics and diverse manipulation skills inherent in the real world. The human video data can also be further expanded to incorporate large-scale datasets like Ego4D [grauman2024ego4d] and EPIC-KITCHENS [damen2020epic].
@@ 5 Experiments
To evaluate the effectiveness of our proposed InternVLA-A1, we conduct extensive experiments on 12 real-world tasks and a simulation benchmark. We first compare InternVLA-A1 with existing leading VLA models. Following this, we conduct a series of ablation studies.
@@ 5.1 Evaluation Configuration
We benchmark the policies on three physical robot embodiments: Agibot Genie-1, ARX Lift-2, and ARX AC One. These platforms cover diverse bimanual manipulation capabilities and execution characteristics, enabling us to assess real-robot performance across both long-horizon and contact-rich behaviors under consistent sensing and control setups. All evaluations are performed with the same deployment pipeline across the three robots.
@@
Task setup. We design a real-world task suite comprising ten static manipulation tasks and two dynamic task families:
@@
Static manipulation tasks: Zip Bag, Unscrew Cap, Make Sandwich, Operate Oven, Sort Parts, Sort Rubbish, Sweep Trash, Wipe Stain, Place Markpen, Place Flower.
@@
Dynamic manipulation tasks: Express Sorting tasks and In-motion Ingredient Picking tasks.
@@
Overall, the tasks span articulated-object interaction and contact-rich manipulation (e.g., Operate Oven and Unscrew Cap), long-horizon bimanual manipulation (e.g., Zip Bag and Make Sandwich), and, importantly, high-dynamics settings where the scene evolves and the target is moving during execution. The latter requires the policy to reason about near-future scene changes, so that actions can be generated with upcoming dynamics in mind rather than solely based on the current observation.
@@
Evaluation protocol. We report the average success rate over 30 rollouts per task. Specifically, each task is evaluated under 30 predefined settings (e.g., object placements and scene initializations within bounded ranges), with one trial per setting. Results are averaged across all rollouts to summarize performance across diverse conditions.
5.2 Evaluation on the Static Manipulation Tasks
@@
To evaluate the general manipulation capabilities of InternVLA-A1 against leading VLA models, we conducted a comprehensive evaluation across 10 diverse real-world static tasks (Figure˜4).
@@
As shown in Table 4, InternVLA-A1 clearly outperforms prior state-of-the-art models on real-world static manipulation tasks. Notably, our smaller InternVLA-A1 (2B) achieves an average success rate of 64.7%, surpassing the 60.6% average of the larger (3.3B) baseline. This result highlights the effectiveness of our architecture and joint training of simulated trajectories, real-robot demonstrations, and egocentric human videos, enabling a lighter model to outperform a larger counterpart. When scaling to a comparable size, InternVLA-A1 (3B) delivers a clear performance gain, reaching an average success rate of 75.1%, which is +14.5% over (60.6%) and +4.4% over (70.7%). It shows strong advantages on long-horizon tasks such as Make Sandwich (93.3% vs. 73.3% for ) and Operate Oven (86.7% vs. 80.0%), and remains top-tier on Sort Rubbish (97.3%), matching the best baseline. On challenging manipulation tasks, InternVLA-A1 (3B) matches the strongest baseline on Unscrew Cap (66.7%) and Sort Parts (53.3%), while achieving substantial gains on Zip Bag (73.3% vs. 60.0% for , and 40.0% for ). These results confirm that InternVLA-A1 not only excels in high-level task planning but also possesses superior fine-grained control capabilities.
@@
5.3 Evaluation on Dynamic Manipulation Tasks
To assess robustness under environmental dynamics, we evaluate the policies on two challenging real-world dynamic manipulation tasks: Express Sorting and In-motion Ingredient Picking, where targets are moving during execution. The experimental settings for the Express Sorting and In-motion Ingredient Picking tasks are depicted in Figure˜5. Both involve long-horizon operations within dynamic environments. In the Express Sorting Task, the robot must first determine if the package label is facing upwards. If the label is facing downwards, a four-step sequence is initiated: the right arm first executes a “chasing” grasp along the direction of the conveyor movement, followed by flipping the package, after which the left arm performs a “waiting” grasp. Finally, the package is lifted to present the label to the head-mounted camera. Conversely, if the label is facing upwards, the robot skips the flipping sequence and directly proceeds to the final two steps. In the In-motion Ingredient Picking tasks, two robots coordinate to grasp the ingredients required to assemble a beef sandwich, consisting of two slices of bread, a steak, and a piece of lettuce.
As shown in Figure˜6, InternVLA-A1 establishes a clear lead over all baselines, surpassing even the strongest prior VLA model . Notably, the InternVLA-A1 (3B) variant demonstrates the most pronounced advantage. While InternVLA-A1 (2B) already performs strongly on Express Sorting (70.0%), surpassing (53.3%) by +16.7% and substantially outperforming GR00T N1.5 (40.0%) and (36.7%), scaling to 3B pushes performance to 80.0% on Express Sorting and 93.3% on Ingredient Picking. Notably, on Ingredient Picking, a highly dynamic and timing-sensitive setting, InternVLA-A1 (3B) delivers a substantial gain over (93.3% vs. 66.7%, +26.6%), while and GR00T N1.5 remain at 20.0%. Similarly, on Express Sorting it improves over by +26.7% (80.0% vs. 53.3%). Overall, InternVLA-A1 (3B) achieves an average success rate of 86.7% across these two dynamic tasks. These results highlight that InternVLA-A1 is markedly more robust to environmental dynamics and motion-induced distribution shifts, translating foresight-guided decision making into reliable closed-loop control in dynamic scenarios.
@@ 5.4 Evaluation on Simulation Benchmark
We also evaluate InternVLA-A1 on the RoboTwin 2.0 [robotwin2.0] benchmark, covering 50 bimanual tasks under both Easy (clean) and Hard (domain randomized) settings. All models are fine-tuned on a total of 27,500 demonstrations, consisting of 2,500 clean and 25,000 randomized episodes (i.e., 50 and 500 per task), with results averaged over 100 evaluation trials. As shown in Figure˜7, InternVLA-A1(3B) significantly outperforms by margins of 9.4% and 10.1%, and surpasses the strong baseline by 2.6% in both the Easy and Hard settings.